# Locally similar distances and equality of the induced intrinsic distances

### v1

@@ 5 Locally similar distances and composition
In this section, we consider a particular situation where two distances $d_{1}$ and $d_{2}$ are related by $d_{2}=f\circ d_{1}$. We provide a sufficient condition for $d_{2}\stackrel{{\scriptstyle\operatorname{str}}}{{\sim}}d_{1}$ in terms of $f$. More precisely, for convenience in applications, we provide a sufficient condition for $d_{2}\stackrel{{\scriptstyle\operatorname{str}}}{{\sim}}Qd_{1}$, with $0<Q<+\infty$. According to Proposition 4.7, our condition is also sufficient for $d_{2}\cong Qd_{1}$.
@@ 6 Locally similar distances induce the same intrinsic distance
We start with the following elementary lemma which provides a sufficient condition for one interval to be a subset of another. Figure 5 illustrates such intervals.
@@
Then, we make a conclusion about the left extremes: ys<xsy-s<x-s. ∎
@@
We note by $\mathcal{P}([a,b])$ the set of all Riemannian partitions of $[a,b]$:
$\mathcal{P}([a,b])\subseteq\Bigl\{P\subseteq[a,b]\colon\quad P\ \text{is finite}\quad\wedge\quad a,b\in P\Bigr\}.$
Furthermore, let $\mathcal{Q}$ be the set of such Riemannian partitions of $[a,b]$ that the corresponding $R(x)/2$-neighborhoods cover $[a,b]$:
@@
In other words, we consider the list of intervals $((t_{k}-r_{k},t_{k}+r_{k}))_{k=0}^{m}$, and we are going to prove that each two consecutive intervals from this list intersect.
@@
Notice that $t_{k}-r_{k}\in[a,b]$:
@@
Notice that $k-1\in\{1,\ldots,m-1\}$ because $0\leq j<k-1<k\leq m$. Excluding $t_{k-1}$ from $Q$ we get an element of $\mathcal{Q}$ with a lesser number of elements, which again contradicts the minimality of $Q$.
@@
Furthermore, by the continuity of $\eta$, there exists $r>0$ such that if $s\in[a,b]$ and |st|<r|s-t|<r, then $\eta(s)\in W$. Moreover, by the Archimedes–Exodus principle, there exists $q$ in $\mathbb{N}$ such that $1/q<r$. Thus, for each $t$ in $[a,b]$, the following set is nonempty:
@@
Applying Lemma 6.2, we find $m$ in $\mathbb{N}$ and $t_{0},\ldots,t_{m}$ in $[a,b]$ like in Lemma 6.2.
@@
Let $X$ be a set and $d_{1},d_{2}\colon X\to[0,+\infty)$ be distances on $X$ such that $d_{1}\cong d_{2}$. Then, $d_{1}^{\ast}=d_{2}^{\ast}$.
@@
Let $x,y\in X$. Since $d_{1}$ and $d_{2}$ are topologically equivalent, $\Gamma_{d_{1}}(x,y)=\Gamma_{d_{2}}(x,y)$. Let $\gamma\in\Gamma_{d_{1}}(x,y)$ and $\varepsilon>0$. Given a partition $P$ in $\mathcal{P}$, we enumerate its elements in the ascending order: $P=\{p_{0},p_{1},\ldots,p_{N}\}$, $0=p_{0}<p_{1}<\cdots<p_{N}=1$. Then, for each $j$ in $\{1,\ldots,N\}$, we apply Lemma 6.5 to the subpath $\eta_{j}\coloneqq\gamma|_{[p_{j-1},p_{j}]}$. Thereby, we get $m_{j}$ in $\mathbb{N}$ and a list of real numbers $t_{j,0},t_{j,1},\dots,t_{j,m_{j}}$ such that
@@
Examples 2.10, 2.11, and 3.9 show that $d_{1}^{\ast}=d_{2}^{\ast}$ does not imply that $d_{1}\cong d_{2}$.
@@
Let $X$ be a set and $d_{1},d_{2}$ be distances on $X$ such that $d_{1}\cong d_{2}$ and $d_{2}^{\ast}=d_{2}$. Then, $d_{1}^{\ast}=d_{2}$.
In Examples 2.12 and 2.15, the distances have the property $d^{\ast}\cong d$.
@@ 7 Examples related to kernels
The corresponding RKHS was described by Steinwart, Hush, and Scovel [17]. In this example,
@@
$K(w,z)=\operatorname{e}^{w\,\overline{z}}L_{m-1}(|w-z|^{2}).$ (7.5)
Here $L_{m-1}$ denotes the Laguerre polynomial of degree m1m-1. See also Youssfi [19] for another proof of (7.5), in the multidimensional case. The associated distance is
$d_{K}(w,z)=\sqrt{1-\operatorname{e}^{-|w-z|^{2}}L_{m-1}(|w-z|^{2})^{2}}\qquad(w,z\in\mathbb{C}).$
@@
$f(t)\coloneqq\sqrt{1-\operatorname{e}^{-t^{2}}L_{m-1}(t^{2})^{2}}.$
An elementary analysis shows that $f$ satisfies conditions of Proposition 5.2, with $Q=\sqrt{m}$. Therefore, $d_{K}\cong\sqrt{m}\,d_{\mathbb{C}}$. By Theorem 6.6 and Example 2.9, $d_{K}^{\ast}=\sqrt{m}\,d_{\mathbb{C}}^{\ast}=\sqrt{m}\,d_{\mathbb{C}}$. ∎

### v2

@@ 4 Locally similar distances and composition
In this section, we consider a particular situation where two distances $d_{1}$ and $d_{2}$ are related by $d_{2}=f\circ d_{1}$. We provide a sufficient condition for $d_{2}\stackrel{{\scriptstyle\operatorname{str}}}{{\sim}}d_{1}$ in terms of $f$. More precisely, for convenience in applications, we provide a sufficient condition for $d_{2}\stackrel{{\scriptstyle\operatorname{str}}}{{\sim}}Qd_{1}$, with $0<Q<+\infty$. According to Propositions 3.8 and 3.9, our condition is also sufficient for $d_{2}\cong Qd_{1}$.
@@
5 The intrinsic distance induced by a given distance
In this section, we recall the definition of the intrinsic distance $d^{\ast}$ induced by a given distance $d$. It is a standard construction in the theory of length spaces (see, for example, Gromov [11, Example 1.4 (a)], Burago, Burago, and Ivanov [8, Section 2.3], Kunzinger and Steinbauer [12, Section 1.2.3]). In this section, we also compute $d^{\ast}$ for several examples. In particular, we show that $\stackrel{{\scriptstyle\operatorname{Lip}}}{{\sim}}$ does not imply $\stackrel{{\scriptstyle\operatorname{intr}}}{{\sim}}$ and $\stackrel{{\scriptstyle\operatorname{intr}}}{{\sim}}$ does not imply $\stackrel{{\scriptstyle\operatorname{top}}}{{\sim}}$.
We now fix some notation for the intermediate concepts which will be used in the examples and the proof of the main result.
Definition 5.1 (continuous paths between two points).
Let $(X,d)$ be a metric space. For each $x,y$ in $X$, define
Definition 5.2 (Riemannian partitions of an interval).
Let $a,b]\subseteq\mathbb{R}$ be an arbitrary interval. We denote by $\mathcal{P}([a,b])$ the set of all finite subsets of $[a,b]$ that include $a$ and $b$. In particular, we denote $\mathcal{P}([0,1])$ by $\mathcal{P}$.
Given $P$ in $\mathcal{P}([a,b])$, we enumerate its elements in ascending order:
$P=\{t_{0},t_{1},\ldots,t_{m}\},\qquad a=t_{0}<t_{1}<\cdots<t_{m}=b.$ (5.1)
Lemma 5.4.
$S(d,\gamma,P)\leq S(d,\gamma,Q).$
Proof.
It follows by an iterative application of the triangle inequality. ∎
Definition 5.5 (the length of a path with respect to a distance).
$L(d,\gamma)\coloneqq\sup_{P\in\mathcal{P}}S(d,\gamma,P).$
Definition 5.6 (the intrinsic distance induced by a distance).
$d^{\ast}(x,y)\coloneqq\inf_{\gamma\in\Gamma_{d}(x,y)}L(d,\gamma).$
Following Definition 5.6, if $\Gamma_{d}(x,y)=\emptyset$ (i.e., if there are no continuous paths between $x$ and $y$), then $d^{\ast}(x,y)=+\infty$.
It is well known that $d^{\ast}$ is a $0,+\infty]-$valued distance on $X$, such that $d^{\ast}(x,y)\geq d(x,y)$ for every $x,y$ in $X$. Moreover, it is well known that $(d^{\ast})^{\ast}=d^{\ast}$. These facts are proved in the books cited at the beginning of this section.
Remark 5.7.
In this paper, we define paths as continuous functions with domain $[0,1]$. Similarly, one can consider paths defined on arbitrary closed intervals of the form $[a,b]$. Such generalization does not change $L(d,\gamma)$ or $d^{\ast}(x,y)$, since any path defined on an arbitrary interval $[a,b]$ can be reparametrized to $[0,1]$.
Proposition 5.8.
Proof.
Let $x,y\in X$ and $\gamma\in\Gamma_{d_{1}}(x,y)$. For every $P$ in $\mathcal{P}$,
$S(d_{2},\gamma,P)\leq C_{2}S(d_{1},\gamma,P)\leq C_{2}L(d_{1},\gamma).$
Example 5.9 (Euclidean distance on $\mathbb{R}^{n}$).
Example 5.10 (the intrinsic distance induced by the finite discrete distance).
Let $X$ be a set having more than one point and $d_{0}\colon X^{2}\to[0,+\infty)$ is defined by
$d_{0}(x,y)\coloneqq\begin{cases}0,&x=y;\\ 1,&x\neq y.\end{cases}$
Since $d_{0}$ induces the discrete topology, there are no continuous paths between different points of $X$, and $d_{0}^{\ast}(x,y)=+\infty$ if $x\neq y$. ∎
Example 5.11 (square-root distance on $\mathbb{R}$).
Consider the distance $d_{\operatorname{sqrt}}\colon\mathbb{R}^{2}\to[0,+\infty)$, given by $d_{\operatorname{sqrt}}(x,y)\coloneqq\sqrt{|x-y|}$. The topology induced by $d_{\operatorname{sqrt}}$ is the usual topology in $\mathbb{R}$, since the set of open balls with respect to the Euclidean distance on $\mathbb{R}$ is the same as the set of open balls with respect to the distance $d_{\operatorname{sqrt}}$.
Let $x,y\in\mathbb{R}$. We will show that if $x\neq y$, then $d_{\operatorname{sqrt}}^{\ast}(x,y)=+\infty$. Suppose $x<y$ (the case if $x>y$ is similar). Let $\gamma\in\Gamma_{d_{\operatorname{sqrt}}}(x,y)$. Notice that $\gamma$ is continuous in the usual sense. Let $m\in\mathbb{N}$. For every $j$ in $\{0,\ldots,m\}$, we define
$z_{m,j}\coloneqq x+\frac{j}{m}(y-x).$
$t_{m,m}\coloneqq 1,$
$t_{m,0}<t_{m,1}<\cdots<t_{m,m}.$
Taking the supremum over $m$ in $\mathbb{N}$ shows that $L(d_{\operatorname{sqrt}},\gamma)=+\infty$. Since $\gamma$ was an arbitrary path between $x$ and $y$, it follows from Definition 5.6, that $d_{\operatorname{sqrt}}^{\ast}(x,y)=+\infty$ if $x\neq y$.
Let $d_{0}\colon\mathbb{R}^{2}\to[0,+\infty)$ is defined as in Example 5.10. Although $d_{0}$ and $d_{\operatorname{sqrt}}$ are not topologically equivalent, their corresponding intrinsic distances are, in fact, equal: $d_{0}^{\ast}=d_{\operatorname{sqrt}}^{\ast}$. ∎
In the following example we develop the idea from Example 5.11 in such a manner that the induced intrinsic distance is finite.
Example 5.12 (quadrant of the plane with the usual distance added with the sqrt-angular pseudodistance).
Let $X$ be the closed top right quadrant of the closed unit disk on the complex plane:
$X\coloneqq\bigl\{z\in\mathbb{C}\colon\ |z|\leq 1,\ \operatorname{Re}(z)\geq 0,\ \operatorname{Im}(z)\geq 0\bigr\}.$
Given $z$ in $X$, let $\arg(z)$ be the principal argument of $z$ belonging to $0,\pi/2]$. We define $d\colon X^{2}\to[0,+\infty)$ by
$d(z,w)\coloneqq\begin{cases}|z|+|w|,&zw=0;\\[4.30554pt] |z-w|,&zw\neq 0,\ \arg(z)=\arg(w);\\[4.30554pt] |z-w|+\sqrt{|\arg(z)-\arg(w)|},&zw\neq 0,\ \arg(z)\neq\arg(w);\end{cases}$
Furthermore, we define $\rho\colon X^{2}\to[0,+\infty)$ by
$\rho(z,w)\coloneqq\begin{cases}|z-w|,&zw\neq 0,\ \arg(z)=\arg(w);\\[4.30554pt] |z|+|w|,&\text{otherwise}.\end{cases}$
Both $d$ and $\rho$ are distances on $X$. The topology induced by $d$ coincides with the usual topology of the set $X$ induced by the Euclidean distance. The topology induced by $\rho$ is different. Indeed, if $z\neq 0$ and $0<\delta<\min\{|z|,1-|z|\}$, then the ball $B_{\rho}(z,\delta)$ is in fact the segment $\{\lambda z\colon\ |z|-\delta<\lambda<|z|+\delta\}$.
Given $z,w$ in $X$ such that $z\neq 0$, $w\neq 0$, and $\arg(z)\neq\arg(w)$, and a path $\gamma\in\Gamma_{d}(z,w)$ such that $0\notin\gamma([0,1])$, we apply the reasoning from Example 5.11 to $\arg\circ\gamma$ and conclude that $L(d,\gamma)=+\infty$.
Therefore, if $\eta$ is a path of finite length from $z$ to $w$, then it must pass through 0. Moreover, some subpaths of $\eta$ must be contained in the segments joining $z$ with 0 and 0 with $w$. Figure 4 shows a couple of paths: one has finite length with respect to $d$ and the other has infinite length.
It follows that $d^{\ast}=\rho$. Consequently, $\rho^{\ast}=\rho$. In this example, $d$ is not topologically equivalent to $\rho$, but they induce the same intrinsic distance.
[…]
Figure 4: The set $X$ from Example 5.12, two points belonging to $X$ ($z$ and $w$), a path of finite length between them (blue) and a path of infinite length (red), with respect to $d$.
Example 5.13 (pseudologarithmic distance on the halfline).
Let $X\coloneqq(0,+\infty)$. Define $\rho\colon X^{2}\to[0,+\infty)$,
$\rho(x,y)\coloneqq\frac{2|x-y|}{x+y}.$
We are going to prove that
$\rho^{\ast}(x,y)=|\log(x)-\log(y)|.$ (5.2)
Consider the case where $x<y$ (the case $x>y$ is similar).
1. Let $\xi(t)\coloneqq x+t(y-x)$ be the straight-line path between $x$ and $y$, and let $Q\in\mathcal{P}([0,1])$. For each $m$ in $\mathbb{N}$, we add to $Q$ the points $k/m$, where $k\in\{0,1,\ldots,m\}$, and obtain a refined partition $R_{m}=\{t_{m,0},t_{m,1},\ldots,t_{m,m}\}$. For each $k$ in $\{0,1,\ldots,m\}$, we define $z_{m,k}\coloneqq\xi(t_{m,k})$. This setup leads to an explicit formula for the polygonal length of $\xi$ associated with $R_{m}$:
$S(\rho,\xi,Q)$ $\leq S(\rho,\xi,R_{m})=\sum_{k=0}^{m-1}\rho(\xi(t_{m,k}),\xi(t_{m,k+1}))=\sum_{k=0}^{m-1}\rho(z_{m,k},z_{m,k+1})$
$=\sum_{k=0}^{m-1}\frac{2(z_{m,k+1}-z_{m,k})}{z_{m,k+1}+z_{m,k}}=\sum_{k=0}^{m-1}\frac{z_{m,k+1}-z_{m,k}}{\frac{z_{m,k+1}+z_{m,k}}{2}}.$
The resulting sum is a Riemannian sum for the integral of the function $f(t)=\frac{1}{t}$ over the interval $[x,y]$. The points $z_{m,0},z_{m,1},\ldots,z_{m,m}$ form a Riemannian partition of $[x,y]$, and its mesh tends to 0 as $m\to\infty$:
$\min_{k\in\{0,\ldots,m-1\}}\bigl(z_{m,k+1}-z_{m,k}\bigr)=(y-x)\,\min_{k\in\{0,\ldots,m-1\}}\bigl(t_{m,k+1}-t_{m,k}\bigr)\leq\frac{y-x}{m}.$
Therefore,
$S(\rho,\xi,Q)\leq\lim_{m\to\infty}S(\rho,\xi,R_{m})=\int_{x}^{y}\frac{\mathrm{d}{}t}{t}=\log(y)-\log(x).$
Since $Q$ was an arbitrary element of $\mathcal{P}([0,1])$, we have proven that $L(\rho,\xi)\leq\log(y)-\log(x)$.
2. For an arbitrary $\gamma$ in $\Gamma_{\rho}(x,y)$, we can construct a partition $P_{m}=\{t_{m,0},\ldots,t_{m,m}\}$ with $\gamma(t_{m,k})=x+k(y-x)/m$, as in Example 5.11, and obtain
$L(\rho,\gamma)\geq S(\rho,\gamma,P_{m})=\sum_{k=0}^{m-1}\rho\left(x+k\frac{y-x}{m},x+(k+1)\frac{y-x}{m}\right).$
The right-hand side is also a Riemannian sum for the integral $\int_{x}^{y}\mathrm{d}{}t/t$. Notice that the points $\gamma(t_{m,k})$, where $k\in\{0,\ldots,m\}$, form a Riemannian partition of $[x,y]$, and the mesh of this partition equals (yx)/m(y-x)/m. Passing to the limit as $m\to\infty$ we get
$L(\rho,\gamma)\geq\int_{x}^{y}\frac{\mathrm{d}{}t}{t}=\log(y)-\log(x).$
Thereby, we have proven (5.2). ∎
Example 5.14 (pseudologarithmic distance on a segment).
Let $X\coloneqq[1,2]$. Define $\rho\colon X^{2}\to[0,+\infty)$ as
$\rho(x,y)\coloneqq\frac{2|x-y|}{x+y}.$
Example 5.15 (pseudohyperbolic distance on the halfplane).
Let $\mathbb{H}$ be the upper halfplane of the complex plane:
$\mathbb{H}\coloneqq\{z\in\mathbb{C}\colon\ \operatorname{Im}(z)>0\}.$
Define $\rho_{\mathbb{H}}\colon\mathbb{H}^{2}\to[0,+\infty)$,
$\rho_{\mathbb{H}}(ia,ib)=\frac{|a-b|}{a+b}\qquad(a,b>0).$
Indeed, for $z=x+iy$ and $w=u+iv$,
Example 5.16 (pseudohyperbolic distance on the unit disk).
It is well known (see, e.g., [19, Chapter 33] and [22, Section 1.5]) that $2\rho_{\mathbb{H}}^{\ast}$ is the geodesic distance associated with the Poincaré halfplane metric $\mathrm{d}{}x\,\mathrm{d}{}y/y^{2}$, and $2\rho_{\mathbb{D}}^{\ast}$ is the geodesic distance associated with the Poincaré disc metric $\mathrm{d}{}x\,\mathrm{d}{}y/(1-x^{2}-y^{2})^{2}$. Nevertheless, we emphasize that the intrinsic distances $\rho_{\mathbb{H}}^{\ast}$ and $\rho_{\mathbb{D}}^{\ast}$ in these classical examples can be computed without using Riemannian metrics.
@@ 6 Locally similar distances induce the same intrinsic distance
We start with the following elementary lemma which provides a sufficient condition for one interval to be a subset of another. Figure 5 illustrates such intervals. Lemma 6.1 is useful in the proof of Lemma 6.2.
@@
Then, we make a conclusion about the left extremes: ys<xry-s<x-r. ∎
@@
Let $\mathcal{Q}$ be the subset of $\mathcal{P}([a,b])$ such that the corresponding $R(x)/2$-neighborhoods of its elements cover $[a,b]$:
@@
In other words, we consider the list of intervals $((t_{k}-r_{k},t_{k}+r_{k}))_{k=0}^{m}$, and we are going to prove that each pair of consecutive intervals from this list intersect.
@@
Since $t_{k}-r_{k}\in[a,b]$, the following inequalities hold:
@@
Notice that $k-1\in\{1,\ldots,m-1\}$ because $0\leq j<k-1<k\leq m$. Excluding $t_{k-1}$ from $Q$ we get an element of $\mathcal{Q}$ with a smaller number of elements, which again contradicts the minimality of $Q$.
@@
Furthermore, by the continuity of $\eta$, there exists $r>0$ such that if $s\in[a,b]$ and |st|<r|s-t|<r, then $\eta(s)\in W$. Moreover, by the Archimedes–Eudoxus principle, there exists $q$ in $\mathbb{N}$ such that $1/q<r$. Thus, for each $t$ in $[a,b]$, the following set is nonempty:
@@
Applying Lemma 6.2, we find $m$ in $\mathbb{N}$ and $t_{0},\ldots,t_{m}$ in $[a,b]$ satisfying (6.1).
@@
If $X$ is a set and $d_{1},d_{2}\colon X\to[0,+\infty)$ are distances on $X$ such that $d_{1}\cong d_{2}$, then $d_{1}^{\ast}=d_{2}^{\ast}$.
@@
We suppose that $d_{1},d_{2}\colon X\to[0,+\infty)$ such that $d_{1}\cong d_{2}$. Let $x,y\in X$. Since $d_{1}$ and $d_{2}$ are topologically equivalent, $\Gamma_{d_{1}}(x,y)=\Gamma_{d_{2}}(x,y)$. Let $\gamma\in\Gamma_{d_{1}}(x,y)$ and $\varepsilon>0$. Given a partition $P$ in $\mathcal{P}$, we enumerate its elements in the ascending order: $P=\{p_{0},p_{1},\ldots,p_{N}\}$, $0=p_{0}<p_{1}<\cdots<p_{N}=1$. Then, for each $j$ in $\{1,\ldots,N\}$, we apply Lemma 6.5 to the subpath $\eta_{j}\coloneqq\gamma|_{[p_{j-1},p_{j}]}$. Thereby, we get $m_{j}$ in $\mathbb{N}$ and a list of real numbers $t_{j,0},t_{j,1},\dots,t_{j,m_{j}}$ such that
@@
If $X$ is a set and $d_{1},d_{2}$ are distances on $X$ such that $d_{1}\cong d_{2}$ and $d_{2}^{\ast}=d_{2}$, then $d_{1}^{\ast}=d_{2}$.
In Examples 5.13 and 5.16, the distances have the property $d^{\ast}\cong d$.
@@
Notice that $d_{1}\cong d_{2}$ is not a necessary condition for $d_{1}^{\ast}=d_{2}^{\ast}$. We have seen it in Examples 5.11 and 5.12. In that examples, the initial distances where not topologically equivalent.
In the following Example 6.9, two distances $d$ and $\rho$ induce the same topology and the same finite intrinsic distance $d^{\ast}=\rho^{\ast}$, but $d$ and $\rho$ are not locally similar.
Example 6.9.
Let $X$ be the following “comb” in $\mathbb{R}^{2}$; see Figure 6:
Furthermore, $d^{\ast}=\rho$. In other words, we claim that $d^{\ast}(a,b)=\rho(a,b)$ for every $a$, $b$ in $X$. For brevity, we illustrate the idea of the proof for the particular case where $a=(0,0)$ and $b=(1/q,y)$, $q\in\mathbb{N}$, $0<y\leq 1$. If $\gamma\colon[0,1]\to X$ is a continuous path joining $a$ with $b$, then, by connectivity arguments, there exists $u$ in $(0,1)$ such that $\gamma(u)=(1/q,0)$. It is easy to see that $L(d,\gamma|_{[0,u]})\geq 1/q$ and $L(d,\gamma|_{[u,1]})\geq y$; here the domains of the paths are $[0,u]$ and $[u,1]$; see Remark 5.7. Therefore, $L(d,\gamma)\geq\rho(a,b)$ for every $\gamma$ in $\Gamma_{d}(a,b)$, and it is easy to conclude that $d^{\ast}(a,b)=\rho(a,b)$. In particular, $d\leq\rho$.
@@ 7 Examples related to kernels
The corresponding RKHS was described by Saitoh [17] and by Steinwart, Hush, and Scovel [18]. In this example,
@@
$K(w,z)=\operatorname{e}^{w\,\overline{z}}L^{1}_{m-1}(|w-z|^{2}).$ (7.5)
See also Youssfi [20] for another proof of (7.5), in the multidimensional case. Here $L^{1}_{m-1}$ denotes the associated Laguerre(-Sonine) polynomial of degree m1m-1, with parameter $1$:
$L^{1}_{k}(x)=\sum_{j=0}^{k}(-1)^{j}\binom{k+1}{k-j}\frac{x^{j}}{j!}.$ (7.6)
The associated distance is
$d_{K}(w,z)=\sqrt{1-\frac{1}{m^{2}}\operatorname{e}^{-|w-z|^{2}}L^{1}_{m-1}(|w-z|^{2})^{2}}\qquad(w,z\in\mathbb{C}).$ (7.7)
@@
$f(t)\coloneqq\sqrt{1-\frac{1}{m^{2}}\operatorname{e}^{-t^{2}}L^{1}_{m-1}(t^{2})^{2}}.$
It follows from (7.6) that $L_{m-1}^{1}(t^{2})^{2}=m^{2}-m^{2}(m-1)t^{2}+O(t^{4})$. Therefore, an elementary analysis shows that $f$ satisfies conditions of Proposition 4.2, with $Q=\sqrt{m}$. Therefore, $d_{K}\cong\sqrt{m}\,d_{\mathbb{C}}$. By Theorem 6.6 and Example 5.9, $d_{K}^{\ast}=\sqrt{m}\,d_{\mathbb{C}}^{\ast}=\sqrt{m}\,d_{\mathbb{C}}$. ∎